OpenStudy (anonymous):

HEY! CAN YOU PLEASE HELP ME ASAP!! solve for y: 2^(y-3)=1/16

OpenStudy (anonymous):

PLEASE HELP!! I'm stuck :( and i'll fail!

OpenStudy (e.mccormick):

Know how to convert 1/16 into 2^?

OpenStudy (anonymous):

no :(

OpenStudy (e.mccormick):

Well, what are the powers of 2?

OpenStudy (anonymous):

its 2^(y-3)

OpenStudy (e.mccormick):

2^0=1 2^1=2 2^2=4 and so on.

OpenStudy (anonymous):

there are answer choices too: its 1. 1 2. 0 3. -1 4. -2 so whats the answer? please tell me!

OpenStudy (e.mccormick):

We don't tell people answers here. We help them find them. Any of the powers of 2 come up to 16?

OpenStudy (anonymous):

yea, 2^4

OpenStudy (e.mccormick):

Ok, so if 2^4=16, 2^?=1/16?

OpenStudy (anonymous):

2^-4?

OpenStudy (e.mccormick):

Exactly! Noe replace the right hand side of the original with that and see if it makes more sense.

OpenStudy (anonymous):

oh! so the answer is -1?! since 2^(-1-3)=2^-4 which is 1/16?

OpenStudy (e.mccormick):

Yes. That is how you do these. make the base match, then you can look at the exponents.